OpenStudy (anonymous):

What is the third term in the expansion of (x + 2)^6?

OpenStudy (campbell_st):

do you know about binomial theory..?

OpenStudy (kaylardz0405):

replace x for 3 and solve

OpenStudy (anonymous):

umm not really

OpenStudy (anonymous):

oh ok

OpenStudy (campbell_st):

ok... so can you calculate a combination...?

OpenStudy (anonymous):

so x = the term, right?

OpenStudy (kaylardz0405):

yes

OpenStudy (anonymous):

so the answer is 15625?

OpenStudy (kaylardz0405):

correct

OpenStudy (anonymous):

Yay! Thank you both so much!

OpenStudy (kaylardz0405):

your welcome

OpenStudy (campbell_st):

ok... so this question uses combinations to calculate part of the coefficient of the term in a degree 6 expansion you will end up with 7 terms... the terms start \[t_{0}, t{1}, t_{2}.....,t_{6}\] so the 3rd term will have r = 0 then the term is found using \[^6C_{2} (x)^{6 - 2} \times (2)^2\] simplify it and you will have the 3rd term

OpenStudy (campbell_st):

oops should read the 3rd term has r = 2 so then the term is \[^6C_{2} \times x^{6 -2} \times 2^2\]

OpenStudy (anonymous):

so KaylaRdz0405 is wrong?

OpenStudy (campbell_st):

it follows pascals' triangle